# FED: Fast and Efficient Dataset Deduplication Framework with GPU Acceleration

### v1

@@ Abstract
Dataset deduplication plays a crucial role in enhancing data quality, ultimately improving training performance and efficiency of LLMs. A commonly used method for data deduplication is the MinHash LSH algorithm. Recently, NVIDIA introduced a GPU-based MinHash LSH deduplication method, but it remains suboptimal, leaving room for further improvement in processing efficiency. This paper proposes a GPU-accelerated deduplication framework FED that optimizes MinHash LSH for GPU clusters and leverages computationally efficient and partially reusable non-cryptographic hash functions. FED significantly outperforms the CPU-based deduplication tool included in SlimPajama by up to 58.3 times and the GPU-based deduplication tool included in NVIDIA NeMo Curator by up to 8.6 times when processing 1 million documents with a node of four GPUs. Deduplication of 1.2 trillion tokens is completed in just 5.1 hours in a four-node, 16-GPU environment. The related code is publicly available on GitHub (https://github.com/mcrl/FED).
@@ 1 Introduction
Pretrained language models (PLMs) Vaswani et al. (2017) have demonstrated exceptional performance, enabling unprecedented various application advancements. It is well known that the larger the dataset, the better the PLM’s performance Hoffmann et al. (2022).
@@
High-quality datasets are crucial for LLM training because they directly affect the language models’ performance, generalizability, and robustness Gunasekar et al. (2023); Li et al. (2023). A critical aspect of ensuring data quality is the elimination of redundant data entries, a process known as dataset deduplication. It is natural to consider that duplicated data in the training dataset can make training an LLM more time-consuming and resource-intensive. It may also introduce bias into the model’s learning process towards near-duplicates in the dataset Lee et al. (2022).
@@
Duplicate detection methods can be broadly categorized into two approaches: exact matching, which identifies identical text, and approximate matching, which identifies documents with similar content. Approximate matching is significantly more computationally expensive and time-consuming Albalak et al. (2024). One commonly used approximate matching method is MinHash LSH Indyk and Motwani (1998), a simplified version of MinHash Broder (1997). However, widely adopted implementations of MinHash LSH are too slow to handle large-scale datasets effectively.
Recently, NVIDIA introduced NeMo-Curator Jennings et al. , a data processing toolkit that includes a GPU-accelerated deduplication tool. While it is faster than traditional CPU-based implementations, processing datasets with trillions of tokens still takes over a day. This fact emphasizes the need for further innovation in deduplication techniques to address the ever-increasing scale of PLM datasets.
@@
This paper presents a GPU-based deduplication framework called FED, which significantly outperforms the existing deduplication tools. By leveraging a simple yet efficient hashing function and optimizing GPU kernel operations and communication, we have developed a framework capable of deduplicating a dataset of 1.2 trillion tokens in just 5.1 hours. The contributions of this paper are summarized as follows:
  • GPU-accelerated deduplication. We present a new GPU-accelerated deduplication framework, called FED, optimized for computation and communication. We successfully deduplicated a dataset with one trillion tokens in about five hours.
@@
  • Computationally efficient hash functions. We introduce novel non-cryptographic hash functions that reduce computational cost by partially reusing the previous result during deduplication. As a result, FED can process the hash generation part more than 1,700 times faster than the existing CPU baseline and over 100 times faster than the GPU baseline.
@@
  • Stable deduplication quality. By comparing the deduplication results of the widely used MinHash implementation with our method, we have confirmed that FED provides both speed and reliability in deduplication.
@@ 2.1 Effects of Data Deduplication
Data deduplication involves identifying and removing redundant data to ensure each instance is unique, enhancing learning efficiency and improving data quality. Figure 1 illustrates duplicate data within the RealNews Zellers et al. (2019) dataset. The sentence structures of the two documents are nearly identical, and such redundancy can hinder effective learning for language models (LMs). Lee et al. (2022) suggests that dataset deduplication positively impacts the validation perplexity of language models, and Allamanis (2019) presents that duplicated data in code datasets negatively affects LMs on code understanding tasks. Moreover, train-test leakage, where test data is inadvertently included in the training set, can significantly skew model evaluation metrics. For example, near-duplicates in both training and test sets can lead to overestimating model performance. Thus, deduplication helps provide a more accurate assessment of model capabilities by eliminating these overlaps Lee et al. (2022).
Training on deduplicated datasets also reduces computational resource requirements, as the model processes unique information only once. This results in faster training times and lower costs. Deduplication also mitigates the risk of overfitting by preventing the model from memorizing repeated data entries, leading to better generalization of unseen data. Lee et al. (2022) states that deduplication also helps reduce some privacy concerns as it reduces the likelihood of models generating memorized content from training data, which could include sensitive information.
@@ 2.2 MinHash
MinHash Broder (1997) is a technique approximating the Jaccard similarity (please see Appendix A) between two documents. Identifying similar (duplicate) sets using MinHash can be broken down into four steps: MinHash generation, duplicate pair generation, constructing a union graph, and generating the final list of duplicates.
@@
First, each document is broken down into smaller components, known as shingles or n-grams and represented by the set of its shingles. Then, the shingles are mapped to zero or positive integers using $H$ hash functions, $f_{1},f_{2},\cdots,f_{H}$. As a result, we have $H$ sequences of integers for a document. A MinHash signature for the document is a sequence of the minimum values obtained from each hash function, resulting in a sequence of length $H$. The Jaccard similarity between the two documents is then estimated by comparing their MinHash signatures. $H$ different hash functions result in a $H\times N$ signature matrix for $N$ documents. The choice of hash functions and the number of hash functions are critical hyperparameters in document deduplication based on MinHash. Please, see Appendix B for the details of MinHash generation.
@@
Specifically, MinHash approximates the Jaccard similarity by counting the same elements in the two signatures element-wise:
@@ 2.3 MinHash LSH
The critical difference between MinHash and MinHash LSH lies in the stage of generating duplicate pairs. In MinHash LSH, the signature column vector of each document is divided into $b$ bands, each of which has $r$ integers. LSH hashes the bands in each document to $k$ buckets. When two documents have bands hashed to the same bucket, they are potentially similar documents. Please, see Appendix C for the details of MinHash LSH.
MinHash LSH is the most commonly used method for document deduplication Albalak et al. (2024). Brown et al. (2020) use MinHash LSH with ten hash functions to filter the training dataset for GPT-3. Rae et al. (2021) use MinHash LSH with 13-grams and 450 hash functions for the gopher training dataset.
@@ 2.4 Existing MinHash LSH Implementations
The MinHash-LSH algorithm, implemented using the Python library datasketch, is widely used in many NLP studies. SlimPajama Shen et al. (2023) employed CPU-based MinHash LSH using this library. We identified and fixed some bugs in their implementation, and hereafter, we will refer to this corrected version as the CPU baseline throughout the paper. The CPU baseline operates identically to the MinHash LSH process described in Section 2.3 and Appendix C, with one key difference.
The difference lies in how comparisons are performed between documents grouped into the same bucket. When a bucket contains $B$ documents, comparisons are performed with a $O(B^{2})$ time complexity. However, in the CPU baseline, comparisons are conducted between the first document’s signature vector entering the bucket and those of subsequently arriving documents, resulting in $O(B)$.
Recently, NVIDIA introduced a framework called NeMo Curator Jennings et al. , which includes a GPU-accelerated MinHash LSH algorithm using cuDF RAPIDS and Dask libraries Dask . We will refer to it as a GPU baseline. It is based on the approach used for Megatron-Turing NLG Smith et al. (2022). Similar to the CPU baseline, it does not compare all pairs of documents within a bucket containing $B$ documents. An anchor document for each bucket is selected, and the similarity between it and other documents in the bucket is computed. The GPU baseline is implemented using the Python GPU Data Frame library, cuDF, where document data is converted into a data frame format and processed with GPU-accelerated operations.
@@ 3 Design and Implementation of FED
We present a deduplication framework optimized for GPU cluster environments where each node has multiple GPUs. Considering scenarios where the size of the document set (e.g., trillions of tokens) exceeds the CPU and GPU memory capacity of the cluster, we adopt a strategy to store intermediate results, such as the signature matrix and bucket IDs, on storage devices during the deduplication process.
3.1 Minhash LSH
Assume that the dataset consists of multiple files and that they are stored in a suitable format, such as JSONL. Also, assume that each node in the cluster has $N_{\mbox{GPU}}$ processes when $N_{\mbox{GPU}}$ is the number of GPUs in each node. Each process reads and processes one file at a time. The file is loaded into the CPU’s main memory using two buffers according to a predefined format: one stores the text of all the documents in the file, and the other stores the document indices. These buffers are then transferred to the GPU. On the GPU, multiple CUDA threads generate the MinHash signature matrices for the documents in parallel and create bucket IDs for each band. The signature matrices and bucket IDs are transferred back to the CPU’s main memory and saved to a file. Consequently, each input file has a corresponding hash result file.
In MinHash LSH, documents sharing the same bucket ID are grouped together, and pairwise comparisons of their signature vectors are performed. Since documents with the same bucket ID are distributed across all hash result files, it requires reading all the hash result files, introducing a file I/O bottleneck. To address this problem, we store the documents with the same bucket ID in the same buffer and transfer the buffer to the GPU. Then, the GPU compares the document pairwise.
Unlike the CPU or GPU baseline implementation, which selects a representative document within a bucket containing $B$ documents and performs comparisons in $O(B)$, FED compares all pairs of documents in the bucket in $O(B^{2})$. Despite the increased number of comparisons, FED’s optimized GPU comparison kernel outperforms existing methods. By performing more comparisons, we aim to reduce false negatives.
The similarity computation results are then transferred back to the CPU, where similar documents are connected to build union graphs. Finally, these graphs are merged to group duplicate documents.
@@ 3.2.1 Hashing in the CPU baseline
where $H$ is the number of different hash functions, and $u$ refers to shingles composed of $n$-grams. CPU baseline uses a commonly used secure hash function SHA-1 Eastlake 3rd and Jones (2001), which is relatively slow when processing large data.
LSH divides the signature vector of each document into $b$ bands and hashes the values of each band into $k$ buckets. In the CPU baseline, documents are mapped to the same bucket only if their signature vectors are identical in a particular band. As a result, the bucket size, $k$, tends to be very large in the CPU baseline.
3.2.2 Hashing in the GPU Baseline
The primary focus of MinHash generation is preventing collisions rather than cryptographic security. However, the CPU baseline uses SHA-1, which guarantees cryptographic security. The GPU baseline uses the non-cryptographic MurmurHash3 algorithm for MinHash generation. It produces a 32-bit or 128-bit hash value, enabling faster computation compared to SHA-1.
In the LSH stage, the values of each band are mapped to buckets using the results of the MD5 hash function, which generates a 128-bit fingerprint by encoding a string of any length.
3.2.3 Hashing in FED
@@
Instead of calculating $f(s)$ for each shingle, only a few operations – a multiplication, two additions, and a single division – are necessary when computing the hash value of the next shingle with that of the current shingle. This significantly reduces the computational cost. Moreover, unlike SHA-1, FED’s hash function is not constrained by the length of the string, making it more efficient for longer strings. We vary $(p,q)$ pairs to generate multiple hash functions simultaneously, supporting parallel computation. The hash function is suitable for MinHash with respect to the four requirements of good hash functions: determinacy, uniformity, collision resistance, and security. Security is not a concern because the hash function is not used for security purposes. Appendix D discusses the suitability of the hash functions for MinHash.
@@
For hashing to generate bucket IDs, FED sums the $r$ values in each band and uses the remainder of the division by the predefined constant number of buckets, $K$. This effectively maps each band, which is defined in the space $\mathbb{N}^{r}$, into the space $\mathbb{N}^{K}$. As $K$ increases, the process of scanning through files to group documents with the same key becomes more frequent. The optimal number of buckets can be expressed as $K=kN^{1/2}$, where $k$ is a constant determined empirically (our experimental result shows that setting $k$ to a value between 1 and 2 is appropriate). Using this relationship, we identify the optimal value of $K$. Appendix E describes the details of obtaining $K$.
4 Optimizations and GPU Acceleration
Initial Processing by the CPU.
As mentioned in Section 3.1, each node in the target cluster has $N_{\mbox{GPU}}$ processes when $N_{\mbox{GPU}}$ is the number of GPUs in each node. The CPU processes handle different JSONL files in parallel. Each JSONL file is loaded into the CPU’s main memory using two buffers: one stores the text of all the documents in the file, and the other stores the document indices. After the buffers are filled with a batch of files, the buffer is transferred to the corresponding GPU for shingling and MinHash generation.
Shingling and MinHash generation on the GPU.
The GPU performs shingling and generates MinHash values for the documents contained in the buffer. Each JSONL file consists of multiple documents, and each SM (Streaming Multiprocessor) within the GPU processes different documents independently. That is, a thread block processes a document, with each thread handling different MinHash functions.
Communication-computation overlapping.
To maximize performance, we implement communication-computation overlapping techniques, such as double buffering. When the GPU performs computation, a buffer containing a batch of files is simultaneously transferred to the GPU.
Hash parameters for GPU compatibility.
The values for $p$ and $q$ described in Section 3.2.3 are carefully selected to address the limited support of the GPU’s 64-bit integer operations, ensuring that the result of the operation remains within the 32-bit integer range.
Storing back hash results and bucket IDs.
Before the pairwise comparison step of signatures, we transfer the hash results and bucket IDs to the CPU. If the CPU memory is insufficient, we store them in secondary storage, such as SSDs and HDDs, to address the issues in the target environment with a limited main memory space.
File I/O operation optimization.
To perform pairwise comparisons, we need to gather all documents that belong to the same bucket. That is, we must search the stored files to locate all documents that belong to a given bucket. Since performing file I/O for each bucket ID would be time-consuming, we search multiple bucket IDs within a single file I/O operation, resulting in a significant file I/O time reduction. The number of bucket IDs processed per file I/O operation is a tunable parameter. However, finding a suitable value that balances the CPU memory limits while sufficiently reducing the file I/O time is crucial.
GPU-based pairwise comparison.

### v2

@@ Abstract
Dataset deduplication plays a crucial role in enhancing data quality, ultimately improving the training performance and efficiency of large language models. A commonly used method for data deduplication is the MinHash LSH algorithm. Recently, NVIDIA introduced a GPU-based MinHash LSH deduplication method, but it remains suboptimal, leaving room for further improvement in processing efficiency. This paper proposes a GPU-accelerated deduplication framework, FED, that optimizes MinHash LSH for GPU clusters and leverages computationally efficient, partially reusable non-cryptographic hash functions. FED significantly outperforms the CPU-based deduplication tool in SlimPajama (using 64 logical CPU cores) by up to 107.2 times and the GPU-based tool in NVIDIA NeMo Curator by up to 6.3 times when processing 30 million documents on a node with four GPUs. Notably, our method dramatically accelerates the previously time-consuming MinHash signature generation phase, achieving speed-ups of up to 260 compared to the CPU baseline. Despite these gains in efficiency, FED maintains high deduplication quality, with the duplicate document sets reaching a Jaccard similarity of over 0.96 compared to those identified by the standard MinHash algorithm. In large-scale experiments, the deduplication of 1.2 trillion tokens is completed in just 6 hours in a four-node, 16-GPU environment. The related code is publicly available on GitHub (https://github.com/mcrl/FED).
@@ 1 Introduction
Pretrained language models (PLMs) Vaswani et al. (2017); Raffel et al. (2020a); Brown et al. (2020a); Touvron et al. (2023a, b); Dubey et al. (2024) have demonstrated exceptional performance, enabling unprecedented various application advancements. It is well known that the larger the dataset, the better the PLM’s performance Hoffmann et al. (2022); Rae et al. (2021a).
@@
High-quality datasets are crucial for LLM training because they directly affect the language models’ performance, generalizability, and robustness Gunasekar et al. (2023); Li et al. (2023). A critical aspect of ensuring data quality is the elimination of redundant data entries, a process known as dataset deduplication. Since datasets derived from internet dumps often contain many duplicate or near-duplicate documents Elazar et al. (2024); Magnusson et al. (2023), it is important to consider that this redundancy in the training dataset can increase the time and resources required for LLM training Sorscher et al. (2022). It may also introduce bias into the model’s learning process towards near-duplicates in the dataset Lee et al. (2022).
@@
Duplicate detection methods can be broadly categorized into two approaches: exact matching, which identifies cases where the text is completely identical or where the hash results of the text are the same, and approximate matching, which identifies documents with similar contents. Approximate matching is significantly more computationally expensive and time-consuming Albalak et al. (2024). One commonly used approximate matching method is MinHash LSH Indyk and Motwani (1998), a simplified version of MinHash Broder (1997). However, widely adopted implementations of MinHash LSH are too slow to handle large-scale datasets effectively.
Recently, NVIDIA introduced NeMo-Curator Jennings et al. (2024), a data processing tool with a GPU-accelerated deduplication tool. Although it is faster than traditional CPU-based implementations, tests in our hardware environment indicate that further enhancements are required for practical usability. In this context, we propose a GPU-based deduplication framework called FED, which achieves fast processing without relying on high-end computational resources—thereby addressing the ever-growing scale of PLM datasets.
@@
Our approach leverages a simple yet efficient hashing function and optimizes GPU kernel operations and communication. As a result, our framework can deduplicate a dataset of 1.2 trillion tokens in just 6 hours using 16 GPUs, each with 32GB of memory. The contributions of this paper are summarized as follows:
Figure 1: The process of MinHash generation.
  • GPU-accelerated deduplication. We present a new GPU-accelerated deduplication framework, called FED, optimized for computation and communication. Especially, it compares the document signatures in a matrix multiplication-like manner and exploits the existing GPU hardware units well. It also improves file I/O performance by exploiting parallelism to handle large-scale datasets.
@@
  • Computationally efficient hash functions. We introduce novel non-cryptographic hash functions that reduce computational cost by partially reusing the previous result during deduplication. As a result, FED can process the hash value generation part more than 260 times faster than the existing CPU baseline.
@@
  • Stable deduplication quality. By comparing the deduplication results of the widely used MinHash implementation with FED, we show that FED provides both speed and reliability in deduplication.
@@
2 Background and Related Work
This section describes the background and related work to FED.
@@ 2.1 Effects of Data Deduplication
Data deduplication involves identifying and removing redundant data to ensure each instance is unique, enhancing a PLM’s learning efficiency and improving data quality. Figure 2 illustrates duplicate data within the RealNews Zellers et al. (2019) dataset. The sentence structures of the two documents are nearly identical, and such redundancy can hinder the effectiveness of learning for language models (LMs). Lee et al. (2022) suggests that dataset deduplication positively affects the validation perplexity of language models, and Allamanis (2019) presents that duplicated data in code datasets negatively affects LMs on code understanding tasks. Moreover, train-test leakage, where test data is inadvertently included in the training set, can significantly skew model evaluation metrics. For example, near-duplicates in both training and test datasets can lead to overestimating model performance. Thus, deduplication helps provide a more accurate assessment of model capabilities by eliminating these overlaps Lee et al. (2022); Tirumala et al. (2023); Albalak et al. (2024).
Training on deduplicated datasets also reduces computational resource requirements, as the model processes unique information only once. This results in faster training times and lower costsSorscher et al. (2022). Deduplication also mitigates the risk of overfitting by preventing the model from memorizing repeated data entries, leading to better generalization of unseen data.
@@ 2.2 MinHash
MinHash Broder (1997) is a technique approximating the Jaccard similarity between two documents. Jaccard similarity (Jaccard, 1912), $J(A,B)$, is a metric to quantify the similarity of two finite sets, $A$ and $B$:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (1)
Identifying similar (duplicate) sets using MinHash can be broken down into four steps: MinHash generation, duplicate pair generation, constructing a union graph, and generating the final list of duplicates.
@@
Consider Figure 3. First, each document is broken down into smaller components, known as shingles or n-grams and represented by the set of its shingles. Then, the shingles are mapped to zero or positive integers using $H$ hash functions, $f_{1},f_{2},\cdots,f_{H}$. As a result, we have $H$ sequences of integers to represent a document. A MinHash signature for the document is a sequence of the minimum values obtained from each hash function, resulting in a sequence of length $H$. The Jaccard similarity between the two documents is then estimated by comparing their MinHash signatures. $H$ different hash functions result in a $H\times N$ signature matrix for $N$ documents. The choice of hash functions and the number of hash functions are critical hyperparameters in document deduplication based on MinHash.
@@
Specifically, MinHash approximates the Jaccard similarity of two documents by counting the same elements in their two signatures element-wise:
@@ 2.3 MinHash LSH
The critical difference between MinHash and MinHash LSH lies in the stage of generating duplicate pairs. In MinHash LSH, the signature column vector in the signature matrix of each document is divided into $b$ bands, each of which has $r$ integers. As shown in Figure 3, LSH hashes the bands in each document to $k$ buckets. Let $S_{i}$ be the signature vector of the document $D_{i}$. For a given signature vector $S_{1}$ and $S_{2}$, when at least a pair of bands $B_{S_{1}}$ from $S_{1}$ and $B_{S_{2}}$ from $S_{2}$ hashed to the same bucket, we tag them as candidate pairs, potentially similar documents. For example, in Figure 3, $D_{2}$ and $D_{N}$ are potentially similar documents because they have bands hashed to the same bucket. Then, LSH performs pairwise comparisons on each document’s signature vectors of length $b\times r$ for the candidate pairs. If the similarity between the two vectors exceeds a predefined threshold, the two documents are considered a duplicate pair. Once the union graph is constructed for each band, the union graphs for all bands are merged to obtain the final union graph.
MinHash LSH is the most commonly used method for document deduplication Albalak et al. (2024). Brown et al. (2020b) use MinHash LSH with ten hash functions to filter the training dataset for GPT-3. Rae et al. (2021b) use MinHash LSH with 13-grams and 450 hash functions for the gopher training dataset.
@@
(a) Generating the signature matrix and calculating the bucket ID for each band.
(b) Pairwise comparison and the union graph generation.
Figure 4: The overview of FED.
@@ 2.4 Existing MinHash LSH Implementations
The MinHash-LSH algorithm, implemented using the Python library datasketch, is widely used in many NLP studies. SlimPajama Shen et al. (2023) employed CPU-based MinHash LSH using this library. We identified and fixed some bugs in their implementation, and hereafter, we will refer to this corrected version as the CPU baseline throughout the paper. The CPU baseline operates identically to the MinHash LSH process described in Section 2.3, with one key difference. The difference lies in how comparisons are performed between documents grouped into the same bucket. When a bucket contains $B$ documents, comparisons are performed with a $O(B^{2})$ time complexity. However, in the CPU baseline, comparisons are conducted between the first document’s signature vector entering the bucket and those of subsequently arriving documents, resulting in $O(B)$.
Recently, NVIDIA introduced a framework called NeMo Curator Jennings et al. (2024), which includes a GPU-accelerated MinHash LSH algorithm using cuDF RAPIDS (2024) and Dask libraries Dask (2024). We will refer to it as a GPU baseline. It is based on the approach used for Megatron-Turing NLG Smith et al. (2022). Similar to the CPU baseline, it does not compare all pairs of documents within a bucket containing $B$ documents. An anchor document for each bucket is selected, and the similarity between it and other documents in the bucket is computed. The GPU baseline is implemented using the Python GPU Data Frame library, cuDF, where document data is converted into a data frame format and processed with GPU-accelerated operations.
@@ 3 Design and Implementation of FED
We develop a deduplication framework, FED, optimized for GPU cluster environments where each node has multiple GPUs. It is capable of efficiently performing deduplication on massive datasets containing trillions of tokens. Considering scenarios where the size of the document set (e.g., trillions of tokens) exceeds the CPU and GPU memory capacity of the cluster, FED adopts a strategy to store intermediate results, such as the signature matrix and bucket IDs, on storage devices during the deduplication process. The overall process is illustrated in Figure 4.
3.1 Overview of FED’s Minhash LSH
Assume that the raw dataset consists of multiple files and that they are stored in a suitable format, such as JSONL, as shown in Figure 4(a). Also, assume that each node in the cluster has $N_{\mbox{GPU}}$ processes (each running in a CPU core) when $N_{\mbox{GPU}}$ is the number of GPUs in each node. Each process reads and processes one file at a time. The file is loaded into the CPU’s main memory using two buffers according to a predefined format: one stores the text of all the documents in the file, and the other stores the document indices. These buffers are then transferred to the GPU. On the GPU, multiple CUDA Luebke (2008) threads generate the MinHash signature matrices for the documents in parallel and generate bucket ID for each band. The signature matrices and bucket ID matrices are transferred back to the CPU’s main memory and saved to a file. Consequently, each input file has a corresponding result file after hashing.
In MinHash LSH, documents sharing the same bucket ID are grouped together, and pairwise comparisons of their signature vectors are performed. As shown in Figure 4(b), process $j$ scans all the hash result files and is in charge of $B=N_{band}/N_{GPU}$ bands from $(j-1)\times B$ to $j\times B$. For each band, it extracts the document signatures with the same bucket ID and stores them in a separate buffer for the bucket ID. Due to the GPU memory limitation issues, each process handles $C$ buckets at a time when scanning all the files and iterates until it covers all the buckets. The buffers are then transferred to the GPU for pairwise comparisons at each iteration. The similarity computation results are then transferred back to the CPU, where all the similar documents are connected to build the final union graph (by Process 1).
Unlike the CPU or GPU baseline implementation, which selects a representative document within a bucket containing $B$ documents and performs comparisons in $O(B)$, FED compares all pairs of documents in the bucket in $O(B^{2})$. Despite the increased number of comparisons, FED’s optimized GPU comparison kernel outperforms the existing methods. By performing more comparisons, we aim to reduce false negatives.
@@ 3.2 Hash Functions
where $H$ is the number of different hash functions, and $u$ refers to shingles composed of $n$-grams. CPU baseline uses a commonly used secure hash function SHA-1 Eastlake 3rd and Jones (2001), which is relatively slow when processing large data. LSH divides the signature vector of each document into $b$ bands and hashes the values of each band into $k$ buckets. In the CPU baseline, documents are mapped to the same bucket only if their signature vectors are identical in a particular band. As a result, the bucket size, $k$, tends to be very large in the CPU baseline.
Hashing in the GPU Baseline
The primary focus of MinHash generation is preventing collisions rather than cryptographic security. However, the CPU baseline uses SHA-1 Eastlake 3rd and Jones (2001), which guarantees cryptographic security. The GPU baseline uses the non-cryptographic MurmurHash3 algorithm Appleby (2012) for MinHash generation. It produces a 32-bit or 128-bit hash value, enabling faster computation than SHA-1. In the LSH stage, the values of each band are mapped to buckets using the results of the MD5 hash function Rivest (1992), which generates a 128-bit fingerprint by encoding a string of any length.
Hashing in FED
@@
Instead of calculating $f(s)$ for each shingle, only a few operations – a single multiplication, two additions, and a single division – are necessary when computing the hash value of the next shingle with that of the current shingle. This significantly reduces the computational cost. Moreover, unlike SHA-1, FED’s hash function is not constrained by the length of the string, making it more efficient for longer strings. We vary $(p,q)$ pairs to generate multiple hash functions simultaneously, supporting parallel computation.
Number of Buckets in LSH
@@
For hashing to generate bucket IDs, FED sums the $r$ values in each band and uses the remainder of the division by the predefined constant number of buckets, $K$. This effectively maps each band, which is defined in the space $\mathbb{N}^{r}$, into the space $\mathbb{N}^{K}$. As $K$ increases, the process of scanning through files to group documents with the same key becomes more frequent. This process is proportional to $K$, resulting in $O(K)$. On the other hand, the number of documents with the same key is, on average, $O\left(\frac{N}{K}\right)$ when $N$ is the total number of documents. Since the comparison time is proportional to the square of the number of documents, the overall time complexity becomes $O\left(\left(\frac{N}{K}\right)^{2}K\right)=O\left(\frac{N^{2}}{K}\right)$. Thus, the processing time can be approximated as $T(K)=aNK+\frac{bN^{2}}{K}$, where $a$ and $b$ are some constants. To minimize $T(K)$, we differentiate it with respect to $K$ and set $\frac{dT(K)}{dK}=0$, yielding $K^{2}=\frac{bN^{2}}{aN}$. Solving for $K$, we find $K=\sqrt{\frac{b}{a}}N^{1/2}$. Thus, the optimal number of buckets can be expressed as $K=kN^{1/2}$, where $k=\sqrt{\frac{b}{a}}$ is a constant determined empirically. For example, we set $k$ to 2 for 30B-token dataset. As the dataset size increases, it is recommended to use a larger $k$.
Suitability of FED’s Hash Functions
In general, a good hash function for MinHash should have the following properties:
  • Determinacy: The same input produces the same output.
The proposed hash function satisfies these properties. Determinacy and uniformity are straightforward. Collision resistance is also met with a large value of $p$. When uniformity is satisfied, the probability of a collision approaches $\frac{1}{p}$, which is very small. In many contexts, hash functions are required to resist preimage attacks, making it computationally difficult to reverse-engineer the input from the hash value. However, it is not a concern in MinHash because the hash function is not used for security purposes.
3.3 GPU-based Pairwise Comparison